A man died and left the following instructions for his fortune. Half to his wife, one-seventh of what was left to his son, two-thirds of what was left went to his butler, the man’s pet got the remaining 100,000 pesos. How much money did the man leave behind altogether?


1
Expert's answer
2021-10-05T17:55:45-0400

Man leave A money altogether.

To wife: "A*\frac{1}{2}=\frac{A}{2}." Left "1-\frac{A}{2}=\frac{A}{2}."

To son: "\frac{A}{2}*\frac{1}{7}=\frac{A}{14}." Left "\frac{A}{2}-\frac{A}{14}=\frac{3A}{7}."

To butler: "\frac{3A}{7}*\frac{2}{3}=\frac{2}{7}." Left "\frac{3A}{7}-\frac{2A}{7}=\frac{A}{7}."

To pet: "\frac{A}{7}=100000."

Thus, "A=700000".